THE POINT OF INTERSECTION OF THE LINE 2X +3Y = 18
AND THE CIRCLE WHCH HAVE CENTER (0,0) AND RADIUS 5
(A)3,33/11
(B)33/11,3
(C)-3,33/11
(D)-33/11, -3

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
circle with center (0,0) and radius 5 is x2+y2=25..........(1)
given line is 2x+3y=18..............(2)
4x2+4y2=100
2x=18-3y from (2)
4x2=(18-3y)2
100-4y2=(18-3y)2
on solving we get x=3,y=4
Thanks and Regards
Shaik Aasif
askIITians faculty